OpenStudy (anonymous):

PLEASE HELP c: Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form. y = 7x – 10 y = 7x + 10 y = 7x + 18 y = 7x – 18

OpenStudy (anonymous):

Standard form of an equation for a straight line: y=mx+b, m=slope, b=y-intercept .. Slope of given line=∆y/∆x=(4-(-3)/(2-1)=7/1=7 Equation: y=7x+b find b using one of given points (2,4) 4=7*2+b b=4-14=-10 Equation: y=7x-10

OpenStudy (anonymous):

Oh my god ur my savior! c:

OpenStudy (anonymous):

lol (: not a problem!